OpenStudy (anonymous):

The diameter of a circle has endpoints whose coordinates are R(-2, 2) and S(4, 2). Find the equation of the circle. (Hint: Review the formula for the equation of a circle.) (x - ? )^2 + (y - ? )^2 = ?

OpenStudy (anonymous):

i think thats ronq

OpenStudy (anonymous):

(x-4)^2+(y-2)^2=36

OpenStudy (anonymous):

yeah, Amir got it. I was staring at my answer and i'm like.... nooo....

OpenStudy (anonymous):

Amir, isn't it =9 though, and not 36? because 6 is the diameter and isn't it =r^2 which would be 3^2 which would=9

OpenStudy (anonymous):

diameter is 6

OpenStudy (anonymous):

Yeah, and the equation of a circle is (x-H)^2+(x-J)2=r^2 not d^2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

tnx

OpenStudy (anonymous):

so the real answer would be (x-4)^2+(x-2)^2=9

OpenStudy (anonymous):

sure